answersLogoWhite

0

Singular: die

Plural: dice

Nonsense: dices!

The probability is 2/36 = 1/18

User Avatar

Wiki User

∙ 12y ago
User Avatar

What else can I help you with?

Related Questions

What is the probability of not rolling sum of 6 with two fair dice?

The probability of not rolling a sum of six with two fair dice is 1 minus the probability of rolling a sum of six. There are 36 permutations of rolling two dice. Of these, five sum to six, 1+5, 2+4, 3+3, 4+2, and 5+1. The probability, then of rolling a sum of six is 5 in 36. The probability, then of not rolling a sum of six is 31 in 36, or about 0.8611.


What is the probability that you will roll no number greater that 3 with two dices?

The probability of rolling a total of 3 or less on two dice is 3 in 36 or 1 in 12.


What is the probability of rolling a sum of 15 on two fair dice?

Zero is the probability of rolling a sum of 15 on two fair dice; the maximum value is 12.


What is the probability of rolling a sum of 8 when rolling two dice?

5 in 36


What is the probability of rolling a sum of six then a sum of eight?

To find the probability of rolling a sum of six followed by a sum of eight with two six-sided dice, we first calculate each individual probability. The probability of rolling a sum of six is 5/36, and the probability of rolling a sum of eight is 5/36 as well. Since these events are independent, we multiply the probabilities: (5/36) * (5/36) = 25/1296. Thus, the probability of rolling a sum of six followed by a sum of eight is 25/1296.


When rolling two dice what is the probability the sum is greater than 1?

The probability is 1.


What is the probability of rolling two dice and the sum is not divisible by 5?

The probability is 29/36.


What is the probability of rolling a sum of 1 explain?

It depends on "rolling a sum of 1" with what. One die, two dice or more?


What is the probability of rolling a sum of 8 on at least one of two rolls of a pair of number cubes?

The probability of rolling a sum of 8 on one roll of a pair of dice is 5/36.The probability of not rolling a sum of 8 on one roll of a pair of dice is 31/36.The probability of rolling a sum of 8 twice on two rolls of a pair of dice is(5/36)(5/36) = (5/36)2 .The probability of rolling first a sum of 8 and then rolling a sum that is not 8 on thesecond roll is (5/36)(31/36).The probability of rolling a sum that is not 8 on the first roll and rolling a sum of 8in the second roll is (31/36)(5/36).So The probability of rolling a sum of 8 at least one of two rolls of a pair of dice is(5/36)2 + (5/36)(31/38) + (31/36)(5/36) = 0.258487654... ≈ 25.8%.


What is the probability of not rolling a sum of 11 with two fair dice?

Since there are two permutations of two dice that sum to 11, (5-6) and (6-5), and since there are 36 permutations of rolling two dice, the probability of not rolling a sum of 11 is 34 in 36, or 17 in 18, or about 0.9444.


Probability of rolling a pair of dice and getting a total of 12?

to get a 12 you need to roll two a six on each of the two dices. the probability of rolling a six 1/6 on each, and to get six on two, is the 1/6 *1/6 = 1/36


What is the probability of rolling two dice and not getting the sum of two?

35/36

Trending Questions
When two balanced dice are rolled there are 36 possible outcomes. Find the probability that the sum is a multiple of 3 or greater than 4? What are leading questions in an sample survey? Can you be pregnant if you had intercourse took place 20 days after your period ended? What proportion of a normal distribution is located in the tail beyond a z-score if z equals -1.50? How does a researcher determine if a correlation coefficient is significant? You had been having intercourse since 20 days Yes i had been having intercourse since 20 days I want to know whether there is any chance of she becoming pregnantand what are the symptoms pls yes pls m? What is the probability of a down syndrome carrier male and a non-carrier female producing non-carrier children? How many sixths in two thirds? What is the red percent on logo quiz? A certain board game requires the roll of 3 standard dice to determine the outcome of a battle. What is the probability that at least one of the dice will show an even number? One coin is flipped and one eight-sided die is rolled what is the numbers of possible outcomes? How many tons is 8 million g? What are the disadvantages of the experimental method? What is the degree of freedom when the sample size is 150? How many possible outcomes can you get by tossing 5 coins? What does 23.7 days mean? How many 4 diget numbers can you make using 1 2 3 4 with out repeating any of the digits in the same number? What is the difference from having a standard deviation that is a sample and population and how do you tell which one is suited for the question.? Difference between data warehouse and datamart? A large set of data has a mean of 100 and standard deviation of 15Five are addedto each score in data what is the new mean and standard deviation?